# Fixed-domain curve counts for blow-ups of projective space

### v3

@@ 1.4. Acknowledgments
A.C. was supported by SNF-200020-182181. C.L. was supported by an NSF postdoctoral fellowship, grant DMS-2001976, and the MATH+ incubator grant “Tevelev degrees.”
@@ 2.1. Failure of SAE
We give a general construction of $X$ failing SAE.
Proposition 15.
@@
We will prove that for every $g$, there is a divergent sequence $\{n_{k}=n_{k}[g]\}_{k}$ such the general fiber of the map $\tau$ contains maps with singular domain for all $n_{k}$.
@@
where $\mathsf{H}^{\vee}$ is the pullback of the class of a line on $\mathbb{P}^{r}$. Notice that $\beta[d]\cdot K_{X}^{\vee}$ diverges for large $d$. Let $n\geq 0$ be such that equation (1) is satisfied with $\beta=\beta[d]$, i.e.
@@
be the locally closed locus consisting of stable maps where the domain curve consists of a genus $g$ smooth curve mapping to $X$ with class $d\mathsf{H}^{\vee}$ and containing all $n$ of the markings, attached to a smooth rational tail mapping to $X$ with class $rD$. We claim that ${\mathcal{M}}_{\Gamma}$ dominates ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$, so that SAE fails.
To see this, notice that $m=n-D\cdot K_{X}^{\vee}>n$ satisfies the dimensional constraint (1) for $\mathbb{P}^{r}$ with curve class $d\mathsf{H}^{\vee}$. Therefore, since $\mathsf{Tev}^{\mathbb{P}^{r}}_{g,m,dH^{\vee}}>0$ (for example, we have $m\geq d+2$ for large $d$, so we can apply [11, Theorem 1.1]), for general $(C,p_{1},\ldots,p_{m})\in{\mathcal{M}}_{g,m}$ and $x_{1},\ldots,x_{m}\in\mathbb{P}^{r}$ there exists
@@
where $\widetilde{C}$ is obtained from $C$ by attaching a $\mathbb{P}^{1}$ at $p_{n+1}$ and forgetting $p_{i}$ for $i>n+1$. The restriction $\widetilde{f}$ to $C$ is the unique morphism such that $\pi\circ\widetilde{f}=f$, and the restriction to $\mathbb{P}^{1}$ is a fixed degree $r$ cover of $D$. The map $[\widetilde{f}]$ lies in ${\mathcal{M}}_{\Gamma}$, and maps to a general point of ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$, as needed.
@@
Proof of Theorem 8.
Apply Proposition 15 with $D$ equal to the strict transform of the line between any two of the blown up points. $\Diamond$
@@ 2.2. Generalities for Fano varieties
We now collect general statements used to prove SAE for Fano varieties; in this subsection, we will always assume that $X$ is Fano.
Definition 16.
@@
Fix a genus $g\geq 0$ and a non-zero effective curve class $\beta$ on $X$. Assume that
(7) $\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))\leq\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n}).$
(Note: this is weaker than (1).)
@@
be the locally closed locus parametrizing maps whose domain $C$ has the following topological type. $C$ contains a smooth spine component $C_{{\text{sp}}}$ of genus $g$, containing the marked points $p_{n-m+1},\ldots,p_{n}$. Attached to $C_{{\text{sp}}}$, we have trees of rational curves $T^{\prime}_{1},\ldots,T^{\prime}_{a}$, as well as trees of rational curves $T_{1},\ldots,T_{n-m}$, such that the tree $T_{i}$ contains the marked point $p_{i}$ for $i=1,\ldots,n-m$. See also Figure 1. Here $(n-m)+a\geq 0$; we assume further that if $(n-m)+a=0$, that is, if $C$ is irreducible, then (7) is a strict inequality.
@@
Up to permuting indices of the marked points, the ${\mathcal{M}}_{\Gamma}^{(m,a)}$ contain all of the boundary strata that dominate ¯$\overline{\mathcal{M}}_{g,n}$. Therefore, $X$ satisfies SAE if and only if, given $g,n$, where either $g=0$ and $n\geq 1$, or $n$ is large (depending on $X$ and $g$), the space ${\mathcal{M}}_{\Gamma}^{(m,a)}$ fails to dominate ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$ for all $m,a$.
Let $Z$ be an irreducible component of ${\mathcal{M}}_{\Gamma}^{(m,a)}$ and $[f]\in Z$ be a general point. For $n$ large (depending on $X$ and $g$) and for $g=0$ and $n\geq 1$, consider the following two conditions:
@@
    if $a=0$ and the pushforward under $f$ of each component of $T_{i}$ is an ordinary class for $i=1,\ldots,n-m$, then we have the inequality
    $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}$ $]))-n$
    $\leq-(g+1);$
@@
We will prove that if conditions (*) and (**) hold for all $Z$, then $X$ satisfies SAE. This will follow from the next two propositions.
Proposition 17.
Suppose that $a=0$, that, for $i=1,\ldots,n-m$, the pushforward under $f$ of each component of $T_{i}$ is an ordinary class, and that condition (*) holds. Then, $Z$ fails to dominate ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$.
@@
Arguing as in [17, Proof of Proposition 22], we have
$\mathrm{dim}_{[f]}(Z)$ $\leq\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-n$
$+\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n})+m$
where the inequality is strict when $n=m$.
If $m\geq g+1$ and $Z$ dominates ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$ then, by [17, Proposition 13], we have
and this is a contradiction ($Z$ cannot dominate ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$ for dimensional reasons). When instead $m<g+1$, then we conclude using Condition (*). $\Diamond$
Proposition 18.
Suppose that both conditions (*) and (**) hold. Then, $Z$ fails to dominate ${\mathcal{M}}_{g,n}\times X^{n}$.
@@
Because $X$ is Fano, we have
@@
for all $i=1,\ldots,a$. Therefore, arguing as in [17, proof of Proposition 23], we may assume without loss of generality that $a=0$.
If the pushforward under $f$ of each irreducible component of each $T_{i}$ for $i=1,\ldots,s$ is ordinary, then we are done by Proposition 17.
@@
for all $i=1,\ldots,s$. Let $\hat{f}:\hat{C}\rightarrow X$ be the stable map obtained from $f$ by deleting $T_{1},\ldots,T_{s}$ and $x_{1},\ldots,x_{s}$ and let $\widehat{\beta}=\widehat{f}_{*}[\widehat{C}]$. Notice that $[\widehat{f}]$ belongs to an irreducible component $\widehat{Z}$ of
${\mathcal{M}}_{\Gamma}^{(m,0)}\subset\ \overline{\mathcal{M}}_{g,n-s}(X,\widehat{\beta})$
It is enough to show that $\widehat{Z}$ does not dominate ¯$\overline{\mathcal{M}}_{g,n-s}\times X^{n-s}$. This will follow from Proposition 17, once we notice the following two facts. First, from (7) and condition (**) we have
$\widehat{\beta}\cdot K_{X}^{\vee}=\beta\cdot K_{X}^{\vee}-\sum_{i=1}^{s}f_{*}[T_{i}]\cdot K_{X}^{\vee}\leq r(n-s+g-1)-s$
@@
Second, by Equation (7) and Condition (**), we have
$s\leq\frac{\beta\cdot K_{X}^{\vee}}{r+1}=\frac{r}{r+1}\cdot n+\frac{r}{r+1}\cdot(g-1)$
and so for large $n$ also $n-s$ is large and for $g=0$ and $n\geq 1$, also $n-s\geq 1$.
@@
To summarize, we have shown that:
Proposition 19.
Suppose that every irreducible component $Z\subset{\mathcal{M}}_{\Gamma}^{(m,a)}$ as above satisfies Conditions (*) and (**) for large $n$ (depending on $X$ and $g$), and for all $n\geq 1$ when $g=0$. Then, $X$ satisfies SAE.
@@ 2.3. Blow-ups of $\mathbb{P}^{r}$
By [17, Proposition 13], every irreducible component of ¯$\overline{\mathcal{M}}_{0,1}(X,\beta)$ dominating $X$ has dimension equal to $\beta\cdot K_{X}^{\vee}+r-2$, which must be greater than or equal to the dimension of $X$. $\Diamond$
Lemma 22.
@@ 2.4. SAE for ${\mathsf{Bl}}_{q}(\mathbb{P}^{r})$
We use Proposition 19. Let $Z\subseteq{\mathcal{M}}_{\Gamma}^{(m,a)}$ be an irreducible component dominating ¯$\overline{\mathcal{M}}_{g,n}\times X^{n}$ and let $[f]\in Z$ be a general point.
@@ 1. Step 1
    We verify Condition (*). Assume that $k=0$ and that the pushforward under $f$ of each component of $T_{i}$ is an ordinary class for $i=1,\ldots,n-m$. Write $\beta=d\mathsf{H}^{\vee}+k\mathsf{E}^{\vee}$. We distinguish two cases.
@@ 1. Case 1
      Suppose that $f_{*}[C_{{\text{sp}}}]=-k^{\prime}\mathsf{E}^{\vee}$, with $k^{\prime}>0$. Notice that, since the pushforward under $f$ of each component of $T_{i}$ is an ordinary class for $i=1,\ldots,n-m$, by Lemma 20, we must have
@@
      ¯$\overline{\mathcal{M}}_{g,n}(X,-k^{\prime}\mathsf{E}^{\vee})=\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r-1},k^{\prime}),$
      and so by Lemma 22,
      $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}\overline{\mathcal{M}}_{g,n}(X,-k^{\prime}\mathsf{E}^{\vee})$ $\leq\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r-1},k^{\prime}))+rg$
      $=(r-4)(1-g)+k^{\prime}r+n+rg.$
@@
      $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}$ (¯$(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))+(r-1)g$
@@
      $\leq g-1+d-k+(r-1)g$
      $<n-(g+1),$
      where the last equality holds for every $n\geq 1$ when $g=0$ and for $n$ large when $g>0$. Here, in the last two inequalities, we have used (8) and the dimensional constraint (7).

### v4

@@ 1.4. Acknowledgments
A.C. was supported by SNF-200020-182181 and SNF-222363. C.L. was supported by an NSF postdoctoral fellowship, grant DMS-2001976, and the MATH+ incubator grant “Tevelev degrees.”
@@ 2.1. Failure of SAE
We give a general construction of $X$ failing SAE (Definition 1.5).
Proposition 2.1.
@@
We will prove that for every $g$, there is a divergent sequence $\{n_{k}=n_{k}[g]\}_{k}$ and a corresponding sequence of curve classes $\beta_{k}\in H_{2}(X)$ satisfying (1), such that the general fiber of the map $\tau$ contains maps with singular domain for all $n_{k}$.
@@
where $\mathsf{H}^{\vee}$ is the pullback of the class of a line on $\mathbb{P}^{r}$. Notice that $\beta[d]\cdot K_{X}^{\vee}$ diverges for large $d$. Let $n$ be such that equation (1) is satisfied with $\beta=\beta[d]$, i.e.
@@
be the locally closed locus consisting of stable maps where the domain curve consists of a genus $g$ smooth curve mapping to $X$ with class $d\mathsf{H}^{\vee}$ and containing all $n$ of the markings, attached to a smooth rational tail mapping to $X$ with class $rD$. We claim that ${\mathcal{M}}_{\Gamma}$ dominates $\overline{\mathcal{M}}_{g,n}\times X^{n}$, so that SAE fails.
To see this, notice that $m=n-D\cdot K_{X}^{\vee}>n$ satisfies the dimensional constraint (1) for $\mathbb{P}^{r}$ with curve class $d\mathsf{H}^{\vee}$. Therefore, since $\mathsf{Tev}^{\mathbb{P}^{r}}_{g,m,d\mathsf{H}^{\vee}}>0$ (for example, we have $m\geq d+2$ for large $d$, so we can apply [12, Theorem 1.1]), for general $(C,p_{1},\ldots,p_{m})\in{\mathcal{M}}_{g,m}$ and $x_{1},\ldots,x_{m}\in\mathbb{P}^{r}$ there exists
@@
where $\widetilde{C}$ is obtained from $C$ by attaching a $\mathbb{P}^{1}$ at $p_{n+1}$ and forgetting $p_{i}$ for $i>n+1$. The restriction $\widetilde{f}$ to $C$ is the unique morphism such that $\pi\circ\widetilde{f}=f$, and the restriction to $\mathbb{P}^{1}$ is a fixed degree $r$ cover of $D$. The map $\widetilde{f}$] lies in ${\mathcal{M}}_{\Gamma}$, and maps to a general point of $\overline{\mathcal{M}}_{g,n}\times X^{n}$, as needed.
@@
Proof of Theorem 1.8.
Apply Proposition 2.1 with $D$ equal to the strict transform of the line between any two of the blown up points. $\Diamond$
@@ 2.2. Generalities for Fano varieties
We now collect general statements used to prove SAE for Fano varieties; in this subsection, we will always assume that $X$ is Fano. Let $r$ be the dimension of $X$.
Definition 2.2.
@@
Fix a genus $g\geq 0$, a number $n$ of markings satisfying 2g2+n>02g-2+n>0 and a non-zero effective curve class $\beta$ on $X$.
@@
to be the locally closed locus parametrizing maps whose domain $C$ has the following topological type. $C$ contains a smooth spine component $C_{{\text{sp}}}$ of genus $g$, containing the marked points $p_{n-m+1},\ldots,p_{n}$. Attached to $C_{{\text{sp}}}$, we have trees of rational curves $T^{\prime}_{1},\ldots,T^{\prime}_{a}$, as well as trees of rational curves $T_{1},\ldots,T_{n-m}$, such that the tree $T_{i}$ contains the marked point $p_{i}$ for $i=1,\ldots,n-m$. See also Figure 1.
@@
Up to permuting indices of the marked points, the ${\mathcal{M}}_{\Gamma}^{(m,a)}$ contain all of the boundary strata that dominate $\overline{\mathcal{M}}_{g,n}$. Recall that we are interested in the question of whether the general fiber of the map
is contained in ${\mathcal{M}}_{g,n}(X,\beta)$. This amounts to the question of whether ${\mathcal{M}}_{\Gamma}^{(m,a)}$ fails to dominate $\overline{\mathcal{M}}_{g,n}\times X^{n}$ for all $(m,a)\neq(n,0)$.
Let $Z$ be an irreducible component of ${\mathcal{M}}_{\Gamma}^{(m,a)}$ and let $f]\in Z$ be a general point. Consider the following two conditions on $f$.
@@
    $a=0$, and the pushforward under $f$ of each component of $T_{i}$ is an ordinary class for $i=1,\ldots,n-m$. Moreover, we have the inequality
    (7) $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))\leq n-(g+1).$
@@
Proposition 2.3.
Suppose that $g,n,\beta$ satisfy
(8) $\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))\leq\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n}),$
or equivalently that $\beta\cdot K_{X}^{\vee}\leq r(n+g-1)$. Take $a=0$, and let $f\in Z\subset{\mathcal{M}}_{\Gamma}^{(m,0)}$ be as above. Suppose further that if $m=n$ (i.e., $C$ is irreducible), then (8) is a strict inequality.
Assume that condition (*) holds. Then, $Z$ fails to dominate $\overline{\mathcal{M}}_{g,n}\times X^{n}$.
@@
Suppose first that $m<g+1$. We estimate $\mathrm{dim}_{[f]}(Z)$ as in [16, Proof of Proposition 22]. First, the spine curve $C_{{\text{sp}}}$ of $f$ moves in a family of dimension at most n(g+1)n-(g+1) more than the expected, by (7). The remaining components of $C$ are ordinary rational components, which can be attached to $C$ one at a time. Passing through the point of attachment imposes the expected number of conditions on each rational component, so it follows that
$\mathrm{dim}_{[f]}(Z)-\mathrm{vdim}({\mathcal{M}}_{\Gamma}^{(m,a)})\leq n-(g+1)$
as well. On the other hand, we have
$\mathrm{vdim}({\mathcal{M}}_{\Gamma}^{(m,a)})\leq\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))-(n-m)\leq\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n})-(n-m),$
by (8). Therefore,
$\mathrm{dim}_{[f]}(Z)\leq\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n})+m-(g+1)<\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n}),$
so $Z$ cannot dominate $\overline{\mathcal{M}}_{g,n}\times X^{n}$.
Suppose instead that $m\geq g+1$, and that $Z$ dominates $\overline{\mathcal{M}}_{g,n}\times X^{n}$. Then, by [16, Proposition 13], we have
$\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))=\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}])).$
Arguing as in the previous case, it follows after attaching rational components that $\mathrm{dim}(Z)=\mathrm{vdim}({\mathcal{M}}_{\Gamma}^{(m,a)})$. On the other hand, we also have $\mathrm{vdim}({\mathcal{M}}_{\Gamma}^{(m,a)})\leq\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))$ with equality only if $m=n$. Combining with (8), assumed to be strict if $m=n$, we conclude that $\dim(Z)<\mathrm{dim}(\overline{\mathcal{M}}_{g,n}\times X^{n})$, which is a contradiction. $\Diamond$
Proposition 2.4.
Let $X$ be a smooth, projective Fano variety. Assume that, for all $g\geq 0$, there exists an integer $C^{\prime}(X,g)$ with the following property: if $n,\beta$ satisfy $n\geq C^{\prime}(X,g)$ and the inequality (8), then conditions (*) and (**) hold for all components $Z\subset{\mathcal{M}}_{\Gamma}^{(m,a)}\subset{\mathcal{M}}_{g,n}(X,\beta)$. Assume furthermore that we can take $C^{\prime}(X,0)=3$.
Then, $X$ satisfies SAE.
@@
For any $n,\beta$, consider the map
$\tau:\overline{\mathcal{M}}_{g,n}(X,\beta)\rightarrow\overline{\mathcal{M}}_{g,n}\times X^{n}.$
We will show that, if (8) holds, and if in addition
$n\geq\begin{cases}(r+1)(C^{\prime}(X,g)+g)&\text{ if }g\geq 1\\ 3&\text{ if }g\geq 0,\end{cases}$
then every component $Z\subset{\mathcal{M}}_{\Gamma}^{(m,a)}\subset{\mathcal{M}}_{g,n}(X,\beta)$ fails to dominate $\overline{\mathcal{M}}_{g,n}\times X^{n}$, unless $(m,a)=(n,0)$ and the inequality (8) is an equality. This is sufficient to conclude SAE because (8) is weaker than (1).
Fix a component $Z$ and a general point $f\in Z$. Because $X$ is Fano, we have
@@
for all $i=1,\ldots,a$. Therefore, we may assume without loss of generality that $a=0$, by deleting the trees $T_{i}^{\prime}$. The value of $n$ stays the same and $\beta\cdot K_{X}^{\vee}$ decreases strictly, so (8) still holds.
If the pushforward under $f$ of each irreducible component of each $T_{i}$ for $i=1,\ldots,n-m$ is ordinary, then we are done by Proposition 2.3.
@@
for all $i=1,\ldots,s$. Let $\hat{f}:\hat{C}\rightarrow X$ be the stable map obtained from $f$ by deleting $T_{1},\ldots,T_{s}$ and $x_{1},\ldots,x_{s}$ and let $\widehat{\beta}=\widehat{f}_{*}[\widehat{C}$]. Notice that $\widehat{f}$] belongs to an irreducible component $\widehat{Z}$ of
[…]
It is enough to show that $\widehat{Z}$ does not dominate $\overline{\mathcal{M}}_{g,n-s}\times X^{n-s}$, provided that 2g2+(ns)>02g-2+(n-s)>0. This will follow from Proposition 2.3, once we notice the following two facts. First, from condition (**) we have
$\beta\cdot K_{X}^{\vee}-\widehat{\beta}\cdot K_{X}^{\vee}=\sum_{i=1}^{s}f_{*}[T_{i}]\cdot K_{X}^{\vee}\geq(r+1)s,$
@@
$\mathrm{vdim}(\overline{\mathcal{M}}_{g,n-s}(X,\widehat{\beta}))\leq\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))-(r+2)s.$
Combining with (8) yields
@@
If (g,ns)=(0,2)(g,n-s)=(0,2), then the same calculation shows that $\mathrm{vdim}(\overline{\mathcal{M}}_{0,2}(X,\widehat{\beta}))<\dim(X^{2})-1$.
Second, by (8) and condition (**), we have
$s\leq\frac{\beta\cdot K_{X}^{\vee}}{r+1}\leq\frac{r}{r+1}\cdot n+\frac{r}{r+1}\cdot(g-1).$
If $g\geq 1$, then the assumption $n\geq(r+1)(C^{\prime}(X,g)+g)$ gives $n-s\geq C^{\prime}(X,g)$, so we conclude by Proposition 2.3, applied with nsn-s in place of $n$.
If $g=0$, then $n-s\geq\frac{n+r}{r+1}$, hence $n-s\geq 2$. We conclude again by Proposition 2.3 unless ns=2n-s=2. If ns=2n-s=2, then all components (including the spine) of $\hat{C}$ are ordinary, so
$\mathrm{dim}_{\hat{f}}(\overline{\mathcal{M}}_{0,2}(X,\widehat{\beta}))=\mathrm{vdim}(\overline{\mathcal{M}}_{0,2}(X,\widehat{\beta}))<\dim(X^{2}),$
[…]
@@ 2.3. Blow-ups of $\mathbb{P}^{r}$
By [16, Proposition 13], every irreducible component of $\overline{\mathcal{M}}_{0,1}(X,\beta)$ dominating $X$ has dimension equal to $\beta\cdot K_{X}^{\vee}+r-2$, which must be greater than or equal to the dimension of $X$. $\Diamond$
Lemma 2.7.
@@ 2.4. SAE for ${\mathsf{Bl}}_{q}(\mathbb{P}^{r})$
We adopt the setup of Proposition 2.4. The assumption (8) here is
(9) $n\geq\frac{r+1}{r}\cdot d-\frac{r-1}{r}\cdot k-g+1.$
@@ 1. Step 1
    We estimate the quantity
    $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}])),$
    appearing in (7), under the assumptions of (*). There are two cases.
@@ 1. Case 1
      Suppose that $f_{*}[C_{{\text{sp}}}]=-k^{\prime}\mathsf{E}^{\vee}$, with $k^{\prime}>0$. Notice that, since the pushforward under $f$ of each component of $T_{i}$ is an ordinary class for $i=1,\ldots,n-m$, by Lemma 2.5, we must have
@@
      $\overline{\mathcal{M}}_{g,n}(X,-k^{\prime}\mathsf{E}^{\vee})=\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r-1},k^{\prime}),$
      and so by Lemma 2.7,
      $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}\overline{\mathcal{M}}_{g,n}(X,-k^{\prime}\mathsf{E}^{\vee})\leq\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r-1},k^{\prime}))+(r-1)g.\\$
@@
      $\mathrm{dim}_{[f|_{C_{{\text{sp}}}}]}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))$
      $\leq$ $\,\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r-1},k^{\prime}))-\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,f_{*}[C_{{\text{sp}}}]))+(r-1)g$
      $=$ $\,[r(k^{\prime}-g+1)-1]-[(r+1)(-g+1)+k^{\prime}(r-1)-1]+(r-1)g$
@@
      $\leq$ $\,rg-1+d-k.$
      Now, toward condition (*), we wish to verify that, for all sufficiently large $n$, the inequality
      (11) $rg-1+d-k\leq n-(g+1)\Leftrightarrow n-d+k\geq(r+1)g$
      holds. We claim that (11) holds as long as $n\geq(r^{2}+3r+1)g$. Indeed, if $d\geq r(r+2)g$, then by (9), we have
      $n-d+k\geq\frac{d}{r}+\frac{k}{r}-g+1\geq(r+1)g.$
      If $d<r(r+2)g$, then instead we have
      $n-d+k\geq(r^{2}+3r+1)g-(r^{2}+2r)g=(r+1)g.$